OpenStudy (anonymous):

Let f(x)=x+sqrt(1−x) Find the local maximum and minimum values of f using both the first and second derivative tests.

OpenStudy (anonymous):

I'm having a problem trying to solve the first derivative. I get the the point \[-1 equals \frac{ 1 }{ 2\sqrt{1-x} }\] but i don't get how to make it equal 3/4

OpenStudy (anonymous):

please please help

OpenStudy (anonymous):

\[f(x) = x + \sqrt{1 - x}\]\[f'(x) = 1 - \frac{1}{2\sqrt{1 - x}}\]

OpenStudy (anonymous):

f'(x) = 0, then solve for x

OpenStudy (anonymous):

\[1-\frac{ 1 }{ 2\sqrt{1-x} }=0\] \[\frac{ 1 }{ 2\sqrt{1-x} }=1\] \[2\sqrt{1-x}=1\] \[\sqrt{1-x}=\frac{ 1 }{ 2 }\] can you continue??

OpenStudy (anonymous):

no i got there i don't know what to do next

OpenStudy (phi):

You may have made a mistake when you took the derivative of \[ \sqrt{1−x}= (1-x)^{\frac{1}{2}} \] you should get \[ \frac{1}{2} (1-x)^{(\frac{1}{2}-1)} \frac{d}{dx}(-x)\] using the chain rule that simplifies to \[ \frac{1}{2} (1-x)^{-\frac{1}{2}} \cdot -1\] or \[ \frac{-1}{2\sqrt{1-x}} \]

OpenStudy (phi):

so you should solve \[ 1 - \frac{1}{2\sqrt{1-x}} = 0\]

OpenStudy (anonymous):

yes i said that :)

OpenStudy (phi):

to solve, multiply by 2 sort(1-x) to get rid of the denominator move the -1 to the other side square both sides to get rid of the square root.

OpenStudy (anonymous):

thanks i got the right answer

OpenStudy (anonymous):

welcome :)